OpenStudy (anonymous):

-2x + 2y = 2 -4y - 2x = -22 solve by elimination method

OpenStudy (anonymous):

Just like the last problem follow the same steps

OpenStudy (anonymous):

i dont get the steps

OpenStudy (anonymous):

Line up the equations like this -2x + 2y = 2 -2x - 4y =-22 we want to subtract one equation to the other in order to cancel out a variable. It can be either x or y Which one do you want to pick?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

great the first equation has a -2x and so does the second one so in order to cancel the -2x from the second we have to subtract the first equation to the second like this -2x + 2y = 2 -2x - 4y =-22 -2x+2x - 4y-2y =-22-2 and that equals 0 - 6y =-24 simplify to - 6y =-24 Yay now we want to know what y is so we divide both sides by -6. Can you do that from the last equation?

OpenStudy (anonymous):

so i divide -6 into -24 and get 4 so y=4 then i plug 4 in everytime i see a y Right?

OpenStudy (anonymous):

once you know x or y you can just plug that value in into whatever equation you want and solve for the other x or y that you dont have.